# Aligning Multiple Knowledge Graphs in a Single Pass

### v1

@@ 5.7 Efficiency Study
5.8 Pair-wise KG Alignment
@@
TABLE V: Pair-wise KG alignment $Hits@1$ score (%). MultiEA (X) denotes combining our training strategy with the X encoder.
Methods DBP-4 DWY-3
en-zh en-fr en-ja dbp-wiki dbp-yago
MTransE 14.63 15.91 13.54 35.16 44.35
MultiEA (MTransE) 15.29 16.05 15.68 36.05 46.17
Relative Gain $\uparrow$ +4.51 +0.88 +15.81 +2.53 +4.10
IPTransE 19.84 20.95 21.38 43.51 55.19
MultiEA (IPTransE) 20.26 21.01 23.29 43.68 57.06
Relative Gain $\uparrow$ +0.11 +0.33 +8.93 +0.39 +3.38
GCN-Align 21.48 24.01 21.54 44.75 58.63
MultiEA (GCN-Align) 21.57 25.71 22.19 45.48 60.08
Relative Gain $\uparrow$ +0.42 +7.08 +3.02 +1.63 +2.47
KECG 25.71 27.44 26.77 47.15 61.97
MultiEA (KECG) 26.14 29.48 27.49 50.02 62.44
Relative Gain $\uparrow$ +1.67 +7.43 +2.69 +6.09 +0.76
NAEA 24.69 26.92 26.35 46.51 61.37
MultiEA (NAEA) 25.75 28.16 26.95 49.18 61.52
Relative Gain $\uparrow$ +4.29 +4.61 +2.28 +5.74 +0.24
PSR 25.59 28.20 27.87 48.23 60.52
MultiEA (PSR) 26.27 30.18 28.03 51.36 62.14
Relative Gain $\uparrow$ +2.65 +7.02 +0.57 +6.49 +3.12
TABLE VI: Pair-wise KG alignment $Hits@10$ scores (%). MultiEA (X) denotes combining our training strategy with the X encoder.
Methods DBP-4 DWY-3
en-zh en-fr en-ja dbp-wiki dbp-yago
MTransE 45.52 50.13 47.93 60.54 67.09
MultiEA (MTransE) 45.93 52.26 50.19 60.96 67.36
Relative Gain $\uparrow$ +0.90 +4.25 +4.72 +0.69 +0.40
IPTransE 56.23 57.79 60.66 75.96 79.32
MultiEA (IPTransE) 58.52 59.35 62.43 77.16 82.17
Relative Gain $\uparrow$ +4.07 +2.70 +2.92 +1.58 +3.59
GCN-Align 62.03 68.05 63.96 77.58 86.00
MultiEA (GCN-Align) 62.41 70.37 65.22 78.56 88.02
Relative Gain $\uparrow$ +0.61 +3.41 +1.97 +1.26 +2.35
KECG 67.19 71.29 67.51 80.40 88.65
MultiEA (KECG) 68.03 73.10 69.38 80.73 89.78
Relative Gain $\uparrow$ +1.25 +2.54 +2.77 +0.41 +1.27
NAEA 66.41 71.43 67.83 80.67 88.76
MultiEA (NAEA) 68.18 74.17 70.95 81.35 89.37
Relative Gain $\uparrow$ +2.67 +3.84 +4.60 +0.84 +0.69
PSR 67.83 73.04 69.93 82.41 90.26
MultiEA (PSR) 70.04 75.93 71.81 83.97 90.79
Relative Gain $\uparrow$ +3.25 +3.96 +2.69 +1.90 +0.59
5.9 Hyperparameter Study
@@
The sensitivity of $\gamma$ on the two datasets is shown in Figure 6(a) and Figure 6(e), respectively. We can see that on both datasets, our MultiEA achieves the best performance when $\gamma=0.2$. It means that the importance of the first-order similarity is 0.2 and the importance of the other higher-order similarities is 0.8. This indicates that incorporating high-order similarities is very beneficial for the multiple KG alignment task, demonstrating the effectiveness of our proposed inference enhancement module. This finding is also consistent with the observation in Section 5.5.
Second, we study the sensitivity of the margin hyperparameter $\lambda$ introduced in the loss function as described by Eq. (9). The results are shown in Figure 6(b) and Figure 6(f). As we can see, the performance is very poor when $\lambda=0$, because the model cannot separate the positive examples and the negative examples, causing the underfitting issue. There is a clear inflection point when $\lambda=1$, and thus in the other experiments, we set it to 1 by default. After that, the performance gradually declines, probably because the model suffers the overfitting issue when the margin is too large.
Third, we investigate the sensitivity of $\eta$, which has been introduced to describe the groups of negative examples. As shown in Figure 6(c) and Figure 6(g), our model is not sensitive to this hyperparameter. Therefore, in practice, we set $\eta$ to 10 to save computational resources.
Finally, we show the model performance w.r.t. the training ratio in Figure 6(d) and Figure 6(h). As expected, the training ratio shows a clear positive effect on improving the model performance, which is consistent with the findings of many previous studies, e.g., [12, 19, 28, 13, 37, 25].
6 Conclusion
@@ Acknowledgments
This work was supported in part by the National Natural Science Foundation of China under Grants 62133012, 61936006, 62303366, and 62073255, in part by the Innovation Capability Support Program of Shaanxi under Grant 2021TD-05, and in part by the Fundamental Research Funds for the Central Universities under Grants QTZX23039, XJSJ23030.

### v2

@@ V-G Efficiency Study
TABLE V: Pair-wise KG alignment $Hits@1$ score. MultiEA (X) denotes combining our training strategy with the X encoder.
Methods DBP-4 DWY-3
en-zh en-fr en-ja dbp-wiki dbp-yago
MTransE 14.63% 15.91% 13.54% 35.16% 44.35%
MultiEA (MTransE) 15.29% 16.05% 15.68% 36.05% 46.17%
Relative Gain $\uparrow$ +4.51% +0.88% +15.81% +2.53% +4.10%
IPTransE 19.84% 20.95% 21.38% 43.51% 55.19%
MultiEA (IPTransE) 20.26% 21.01% 23.29% 43.68% 57.06%
Relative Gain $\uparrow$ +0.11% +0.33% +8.93% +0.39% +3.38%
GCN-Align 21.48% 24.01% 21.54% 44.75% 58.63%
MultiEA (GCN-Align) 21.57% 25.71% 22.19% 45.48% 60.08%
Relative Gain $\uparrow$ +0.42% +7.08% +3.02% +1.63% +2.47%
KECG 25.71% 27.44% 26.77% 47.15% 61.97%
MultiEA (KECG) 26.14% 29.48% 27.49% 50.02% 62.44%
Relative Gain $\uparrow$ +1.67% +7.43% +2.69% +6.09% +0.76%
NAEA 24.69% 26.92% 26.35% 46.51% 61.37%
MultiEA (NAEA) 25.75% 28.16% 26.95% 49.18% 61.52%
Relative Gain $\uparrow$ +4.29 +4.61 +2.28 +5.74 +0.24
PSR 25.59% 28.20% 27.87% 48.23% 60.52%
MultiEA (PSR) 26.27% 30.18% 28.03% 51.36% 62.14%
Relative Gain $\uparrow$ +2.65% +7.02% +0.57% +6.49% +3.12%
TABLE VI: Pair-wise KG alignment $Hits@10$ scores. MultiEA (X) denotes combining our training strategy with the X encoder.
Methods DBP-4 DWY-3
en-zh en-fr en-ja dbp-wiki dbp-yago
MTransE 45.52% 50.13% 47.93% 60.54% 67.09%
MultiEA (MTransE) 45.93% 52.26% 50.19% 60.96% 67.36%
Relative Gain $\uparrow$ +0.90% +4.25% +4.72% +0.69% +0.40%
IPTransE 56.23% 57.79% 60.66% 75.96% 79.32%
MultiEA (IPTransE) 58.52% 59.35% 62.43% 77.16% 82.17%
Relative Gain $\uparrow$ +4.07% +2.70% +2.92% +1.58% +3.59%
GCN-Align 62.03% 68.05% 63.96% 77.58% 86.00%
MultiEA (GCN-Align) 62.41% 70.37% 65.22% 78.56% 88.02%
Relative Gain $\uparrow$ +0.61% +3.41% +1.97% +1.26% +2.35%
KECG 67.19% 71.29% 67.51% 80.40% 88.65%
MultiEA (KECG) 68.03% 73.10% 69.38% 80.73% 89.78%
Relative Gain $\uparrow$ +1.25% +2.54% +2.77% +0.41% +1.27%
NAEA 66.41 71.43 67.83 80.67 88.76
MultiEA (NAEA) 68.18% 74.17% 70.95% 81.35% 89.37%
Relative Gain $\uparrow$ +2.67% +3.84% +4.60% +0.84% +0.69%
PSR 67.83% 73.04% 69.93% 82.41% 90.26%
MultiEA (PSR) 70.04% 75.93% 71.81% 83.97% 90.79%
Relative Gain $\uparrow$ +3.25% +3.96% +2.69% +1.90% +0.59%
V-H Pair-wise KG Alignment
@@
V-I Hyperparameter Study
@@
The sensitivity of $\gamma$ on the two datasets is shown in Fig. 6(a) and Fig. 6(e), respectively. We can see that on both datasets, our MultiEA achieves the best performance when $\gamma=0.2$. It means that the importance of the first-order similarity is 0.2 and the importance of the other higher-order similarities is 0.8. This indicates that incorporating high-order similarities is very beneficial for the multiple KG alignment task, demonstrating the effectiveness of our proposed inference enhancement module. This finding is also consistent with the observation in Section V-E.
Second, we study the sensitivity of the margin hyperparameter $\lambda$ introduced in the loss function as described by Eq. (9). The results are shown in Fig. 6(b) and Fig. 6(f). As we can see, the performance is very poor when $\lambda=0$, because the model cannot separate the positive examples and the negative examples, causing the underfitting issue. There is a clear inflection point when $\lambda=1$, and thus in the other experiments, we set it to 1 by default. After that, the performance gradually declines, probably because the model suffers the overfitting issue when the margin is too large.
Third, we investigate the sensitivity of $\eta$, which has been introduced to describe the groups of negative examples. As shown in Fig. 6(c) and Fig. 6(g), our model is not sensitive to this hyperparameter. Therefore, in practice, we set $\eta$ to 10 to save computational resources.
Finally, we show the model performance w.r.t. the training ratio in Fig. 6(d) and Fig. 6(h). As expected, the training ratio shows a clear positive effect on improving the model performance, which is consistent with the findings of many previous studies, e.g., [16, 23, 32, 17, 43, 29].
VI Conclusion
@@ Acknowledgments
This work was supported in part by the National Natural Science Foundation of China under Grants 62133012, 61936006, 62425605, and 62303366, and in part by the Key Research and Development Program of Shaanxi under Grant 2024CY2-GJHX-15.